# Efficient onboard multi-task AI architecture based on self-supervised learning

### v1

@@ IV-C Analysis on low-power hardware
Table V reports some results for a $512\times 512\times 4$ input in terms of inference latency and peak power consumption. It should be noticed that average power consumption serves as a validation of whether the method is fully utilizing the available resources, by staying close to the 7W budget, or not. Latency results are averaged over 10 runs with 200 warmup iterations. We compare the proposed multitask architecture with the aforementioned baseline and state-of-the-art architectures for individual tasks. We can notice that the proposed design can solve three tasks (two segmentation (S) and one classification (C)) with a latency that is inferior of several other single-task models. We also notice that INT8 quantization provides almost a factor of 2 speedup. Considering the input resolution has $512\times 512$ spatial locations to be labeled, we can say that the FP32 inference time of 56.77 ms corresponds to a throughput of 4.62 Mpx/s and that the INT8 inference time of 34.67 ms corresponds to a throughput of 7.56 Mpx/s.
@@
Table VI presents a tradeoff analysis for the multitask problem. In this analysis, we investigate what is the total energy consumed to solve the three tasks, as a function of latency and instantaneous power, in relation to the inference quality. The proposed design is compared with different combinations of methods to address the three tasks. These methods need to be run serially as the system is already fully used by each single task. In particular, we select some interesting combinations including fastest baseline (DLMN + DLMN + MN, refer to Table V for acronyms), highest FP32 quality (HR + mViT + MN), highest INT8 quality (HR + HR + ViT), Transformers-only (mViT + mViT + ViT), and CNN-only (DLMN + DLMN + RN50). Taking the proposed design as the reference, we report $E_{\Delta}$ as the percentage difference between the energy in Joules consumed to complete the three tasks, and $Q_{\Delta}$ as the average percentage difference in inference metrics (i.e., the percentage difference in mIOU or F1 is computed for each task and then averaged over the three tasks). It is interesting to notice how the proposed design requires the least energy to complete the tasks. While it does not provide the best quality overall, modest improvements in quality are offset by large increases in energy consumption (e.g., +8% quality requires +374% energy with HR + mViT + MN), highlighting the good tradeoff achieved by the proposed method.
@@
Finally, we report how methods scale as a function of image size in Fig. 4 and Fig. 5. The upper limit in image size dictated by the Jetson 8GB shared system memory is $1024\times 1024$ for all methods except Transformers, which run out of memory for a $1024\times 1024$ input. In general, slightly better efficiency is achieved with a $1024\times 1024$ input reaching 8.91 Mpx/s of INT8 throughput compared to 7.56 Mpx/s for a $512\times 512$ input.
IV-D Backbone Size Ablations
@@
As mentioned above, all experiments conducted so far have used a GhostNetV2 backbone with the $\alpha$ width parameter set to 1.6. Since this parameter influences the number of features in different layers of the network and the input channels for the various heads, we conducted experiments to explore how the network’s parameters, MACs, FLOPs, and performance change with varying $\alpha$ values, reported in Table VII. In this experiment, SSL training for the backbone feature extractor for 100 epochs is followed by supervised training of the segmentation head for the cloud cover problem. It is interesting to notice that the model scales beyond the $\alpha=1.6$ value used in all our experiments with improved mIOU. However, complexity also scales accordingly, so, while we found $\alpha=1.6$ to be a good tradeoff, if a mission desires it can sacrifice some speed for higher quality maintaining the proposed design by choosing $\alpha=2$ or higher.
@@
TABLE VII: Model size ablation for cloud cover segmentation.
@@
IV-E Comparison of Self-Supervised Learning Methods
We conducted a comparative analysis of our local contrastive loss procedure, detailed in Sec. III-B, against the method presented in [45], which was specifically tailored for contrastive learning training on remote sensing RGB images and can be considered a state-of-the-art self-supervised training approach in the remote sensing field.
In order to provide a fair assessment of the training procedure, we used the same GhostNetV2 backbone architecture of our main experiments. The original authors code was adapted to handle transformations over 4 spectral channels. While preserving the original hyperparameters and experimental settings as in the original work, we conducted the pretraining phase on the BigEarthNet dataset. Subsequently, we froze the weights of our GhostNetV2 backbone and conducted a comprehensive evaluation on the cloud cover dataset, replicating the methodology outlined in IV-B. The results are summarized in VIII, and clearly demonstrate the substantial performance enhancement achieved through our self-supervised learning pretraining compared via a combination of local and global contrastive loss, with respect to the methodology proposed in [45].
TABLE VIII: SSL Training Ablation for cloud cover segmentation.

### v2

@@ IV-C Analysis on low-power hardware
Table V reports some results for a $512\times 512\times 4$ input in terms of inference latency, average power consumption and power-normalized latency. The latter is computed as the product between latency and average power normalized by 7W, i.e. the maximum power budget of the system. It should be noticed that average power consumption serves as a validation of whether the method is fully utilizing the available resources, by staying close to the 7W budget, or not. Latency results are averaged over 10 runs with 200 warmup iterations. We compare the proposed multitask architecture with the aforementioned baseline and state-of-the-art architectures for individual tasks. We can notice that the proposed design can solve three tasks (two segmentation (S) and one classification (C)) with a latency that is inferior of several other single-task models. We also notice that INT8 quantization provides almost a factor of 2 speedup. Considering the input resolution has $512\times 512$ spatial locations to be labeled, we can say that the FP32 inference time of 56.77 ms corresponds to a throughput of 4.62 Mpx/s and that the INT8 inference time of 34.67 ms corresponds to a throughput of 7.56 Mpx/s.
@@
Table VI presents a tradeoff analysis for the multitask problem. In this analysis, we investigate what is the total energy consumed to solve the three tasks, as a function of latency and instantaneous power, in relation to the inference quality. The proposed design is compared with different combinations of methods to address the three tasks. These methods need to be run serially as the system is already fully used by each single task. In particular, we select some interesting combinations including fastest baseline (DLMN + DLMN + MN, refer to Table V for acronyms), highest FP32 quality (HR + mViT + MN), highest INT8 quality (HR + HR + ViT), Transformers-only (mViT + mViT + ViT), and CNN-only (DLMN + DLMN + RN50). Taking the proposed design as the reference, we report $E_{\Delta}$ as the percentage difference between the energy in Joules consumed to complete the three tasks, and $Q_{\Delta}$ as the average percentage difference in inference metrics (i.e., the percentage difference in mIOU or F1 is computed for each task and then averaged over the three tasks). We are not surprised that because of the parallel multitask approach, the proposed design requires the least energy to complete the tasks. While it does not provide the best quality overall, modest improvements in quality are offset by large increases in energy consumption (e.g., +8% quality requires +374% energy with HR + mViT + MN), highlighting the good trade-off achieved by the proposed method.
@@
Finally, we present how the methods scale as a function of image size in Fig. 5 and Fig. 6. The upper limit in image size, dictated by the Jetson’s 8GB shared system memory, is $1024\times 1024$ for all methods except Transformers, which run out of memory at this resolution. Generally, slightly better efficiency is achieved with a $1024\times 1024$ input, reaching 8.91 Mpx/s of INT8 throughput compared to 7.56 Mpx/s for a $512\times 512$ input. Indeed, real onboard acquisitions may be significantly larger than $512\times 512$ or $1024\times 1024$. However, a tiling strategy would be adopted onboard, where the large image would be partitioned into tiles as large as the system memory allows. This is also why we present results in terms of throughput, which normalizes latency by image size.
Figure 4: Qualitative comparison of different segmentation models. Top row: cloud cover segmentation; Bottom row: flood segmentation. Left to right: RGB image, Ground Truth, GhostNetv2 + Segmentation Head (SSL), GhostNetv2 + Segmentation Head (SL), DeepLabV3 + MobileNetV3-L, HRNet18, UNet + MobileViT-S.
IV-D Backbone Architecture Ablation
The choice of using GhostNetV2 as the backbone architecture for our main experiments was driven by being the state-of-the-art model among low-complexity architectures. However, one might wonder how a different backbone compares to GhostNetV2 under the specific SSL multitask setting under study. For this purpose, we test MobileNetV3-L as an alternative and present the results in Table VII. It can be noticed that MobileNet offers a different tradeoff between accuracy and latency, being worse on the former and faster for the latter. Whether this is desirable, it depends on the specifics of the mission under design, and, in particular, its speed target.
For a fair comparison, both GhostNetV2 and MobileNetV3-L were trained under identical conditions: the SSL pre-training was conducted for the same number of epochs on the same dataset, using a width multiplier of 1.6 for both networks. Furthermore, the application-specific heads employed for evaluation were consistent with those described in Subsection III-C.
IV-E Backbone Size Ablations
@@
As mentioned above, all experiments conducted on a GhostNetV2 backbone with the $\alpha$ width parameter set to 1.6. Since this parameter influences the number of features in different layers of the network and the input channels for the various heads, we conducted experiments to explore how the network’s parameters, MACs, FLOPs, and performance change with varying $\alpha$ values, reported in Table VIII. In this experiment, SSL training for the backbone feature extractor for 100 epochs is followed by supervised training of the segmentation head for the cloud cover task. It is interesting to notice that the model scales beyond the $\alpha=1.6$ value used in all our experiments with improved mIOU. However, complexity also scales accordingly, so, while we found $\alpha=1.6$ to be a good tradeoff, if a mission desires it can sacrifice some speed for higher quality maintaining the proposed design by choosing $\alpha=2$ or higher.
@@
TABLE VIII: Model size ablation for cloud cover segmentation.
@@
IV-F Comparison of Self-Supervised Learning Methods
We conducted a comparative analysis of our local contrastive loss technique, detailed in Sec. III-B, against the method presented in [50], which was specifically tailored for contrastive learning training on remote sensing RGB images and can be considered a state-of-the-art self-supervised training approach in the remote sensing field.
In order to provide a fair assessment of the training procedure, we used the same GhostNetV2 backbone architecture of our main experiments. The original authors code was adapted to handle transformations over 4 spectral channels, while we preserved the original hyperparameters and experimental settings as in the original work, pretraining the backbone on BigEarthNet dataset. Then, we froze the weights of our GhostNetV2 backbone and conducted a comprehensive evaluation on the cloud cover dataset, replicating the methodology outlined in IV-B. The results are summarized in Table IX and clearly demonstrate the substantial performance enhancement achieved through our self-supervised learning pretraining compared via a combination of local and global contrastive loss, with respect to the methodology proposed in [50].
TABLE IX: SSL Training Ablation for cloud cover segmentation.